Convert between kilodaltons and number of amino acids: 500 amino acids = ____________ kDa 75 kDa...

  1. Convert between kilodaltons and number of amino acids:

500 amino acids = ____________ kDa

75 kDa = _____________ amino acids
Answer #1

The molecular mass of one amino acid is around 110 Da.

So 500 amino acids = 110*500 = 55000 Da or 55kDa

75 kDa = 75000 Da.

So, number of amino acids = 75000/110 = 682 amino acid.

500 amino acids = 55 kDa

75 kDa = 682 amino acids